Growth factors for the population of Chattanooga in the past two years have been 8 and 12. Find the value of the geometric mean.


1
Expert's answer
2021-02-24T07:25:15-0500

The geometric mean is calculated by the formula:


"\sqrt[n]{x_1x_2\cdot\cdot\cdot x_n}"


Growth factors for the population in the past two years have been 8 and 12, therefore the geometric mean is


"\sqrt{12\cdot8}=\sqrt{96}=4\sqrt{6}"

Answer: "4\sqrt{6}"